# Meta-learning of Gibbs states for many-body Hamiltonians with applications to Quantum Boltzmann Machines

### v1

@@ IV.2 Preparing Gibbs state across different regimes of Kitaev ring Hamiltonian
Here, we consider a system size of $L=3$ at temperature of $T=1/\beta=0.1$. The Meta-VQT with $6$ total qubits, $4$ encoding and $1$ processing layer is trained for the subsequent generation of quantum Gibbs state as a function of the strength of transverse field $h$. A total of $20$ $h$ values are chosen uniformly from the interval $[0.7,1.2]$ for training purposes. The reason for this particular choice of interval stems from the fact that as $h$ varies from $0.7$ to $1.2$ the model exhibits transition from semi-classical to quantum critical to semi-classical regime for a temperature value of $T=0.1$, see Fig. 10(b) and (c). For testing the Meta-VQT algorithm, $40$ uniform $h$ values are taken from the interval between $[0.7,1.2]$.
@@ IV.3 Application of NN Meta-VQT: Quantum Boltzman machine
Figure 13: NN Meta-VQT in QBM workflow : (a) and (b) shows the convergence of loss function and the trace distance between exact and prepared Gibbs state using our method of the NN Meta-VQT and existing method of the VarQITE respectively. The KL divergence in (a) converges to $0.0012$ while that in (b) converges to $0.0067$. Additionally, the prepared Gibbs state is found to be more accurate and precise for NN Meta-VQT based QBM than VarQITE based QBM. The number of layers considered for NN Meta-VQT is $4$ SU2 while VarQITE requires $32$ layers of SU2 circuit to prepare Gibbs state with an average value of trace distance $0.22\pm 0.094$.
[…]
@@ V Conclusions
In this work, we introduced two methods of collective optimization known as Meta-VQT and NN Meta-VQT to prepare thermal states of a parametrized Hamiltonian. Unlike the traditional VQT, where the optimization is done for a single parameter configuration of the Hamiltonian, the Meta and NN Meta-VQT is trained to learn the Gibbs state of a set of parameters of the Hamiltonian. This enables efficient generalization across a distribution of Hamiltonians, reducing quantum resource overhead, and allowing rapid inference of thermal states for unseen configurations without retraining from scratch. We accomplish this by encoding the parameters of the Hamiltonian using an encoding layer, which is a SU2 hardware efficient ansatz in Meta-VQT and a neural network in NN Meta-VQT, followed by a processing layer which is HVA in Meta-VQT (Fig. 4) and hardware efficient ansatz and///or HVA in NN Meta-VQT (Fig. 5). The trainable angles in Meta-VQT is provided by the encoding and processing quantum circuits, while for NN Meta-VQT the trainable parameters are in the classical neural network. We observe that the ansatz (Meta-VQT and NN Meta-VQT) trained for a set of $10$ parameters of the Hamiltonian is able to generate Gibbs state close to the exact one corresponding to all other parameters not in the training set, but within the range of parameters chosen for training.
@@ Appendix A Variational Quantum Imaginary Time Evolution (VarQITE)
For an initial state, |$\ket{\psi_{0}}$, the imaginary time $(\tau=it)$ evolution (ITE) with respect to the Hamiltonian, $H(\vec{h})$, non-unitarily maps the state to,
|$\ket{\psi_{\tau}}=\frac{e^{-H(\vec{h})\tau}\ket{\psi_{0}}}{\sqrt{\bra{\psi_{0}}e^{-2H(\vec{h})\tau}\ket{\psi_{0}}}},$ (5)
@@
$\frac{d\ket{\psi_{\tau}}}{d\tau}=-(H(\vec{h})-E_{\tau})\ket{\psi_{\tau}},$
where $E_{\tau}=\bra{\psi_{\tau}}H(\vec{h})\ket{\psi_{\tau}}$ is the system’s energy. If the initial state |$\ket{\psi_{0}}$ is taken to be a purification of the maximally mixed state, $\rho=\text{Tr}_{A}\ket{\psi_{0}}\bra{\psi_{0}}=\mathbb{I}/d$, where A represents the ancilla subsystem, and the imaginary time evolution is kept as $\tau\rightarrow 1/2T=\beta/2$ for the system, then Eq. (5) yields a purification of the Gibbs state defined in Eq. (1).
The ITE state cannot be directly implemented on a quantum circuit because of the non-unitary nature of evolution. However, McArdle et al.[33] showed that the ITE can be implemented on a hybrid gate-based quantum computer using McLachlan’s variational principle. The main idea here is as follows. First, the state |$\ket{\psi_{\tau}}$ is approximated by defining a parametrized trial state : |$\ket{\phi(\vec{\omega}(\tau))}=U(\vec{\omega}(\tau))\ket{0}$, where $U(\vec{\omega}(\tau))=U_{N}(\omega_{n})U_{N-1}(\omega_{N-1}).....U_{k}(\omega_{k})....U_{1}(\omega_{1})$ are the set of parametrized single or two-qubit gates with parameters $\vec{\omega}(\tau)$. Since the goal is the preparation of Gibbs state and this requires purification of maximally mixed states as the initial boundary condition, the state |$\ket{\phi(\vec{\omega}(\tau))}$ is defined over a PQC with a total of $\text{n}+\text{n}_{\text{ancilla}}$ qubits. The initial value of parameters of the unitaries are chosen such that,
|$\ket{\phi(\vec{\omega}(\tau=0))}=U(\vec{\omega}(\tau=0))\ket{0}^{{\otimes}2n}=\ket{\phi^{+}}^{\otimes n},$
where |$\ket{\phi^{+}}^{\otimes n}=\frac{1}{\sqrt{2}}(\ket{00}+\ket{11})$ is the Bell state. Here, in each Bell state |$\ket{\phi^{+}}$, the first qubit belongs to the system, while the second qubit serves as an ancilla. Tracing out the ancillas from |$\ket{\phi(\omega(0))}$ results in the maximally mixed state; $\frac{\mathbb{I}}{2^{n}}$. Further, the ITE of |$\ket{\phi(\vec{\omega}(\tau))}$ is done by implementing McLachlan variational principle,
$\delta||(\frac{\partial}{\partial\tau}+H(\vec{h})-E_{\tau})\ket{\phi({\vec{\omega}(\tau)})}||=0,$
using quantum circuits for $\tau=\beta/2$. The state, $\text{Tr}_{b}[\ket{\phi(\vec{\omega}(\tau=\beta/2))}]$ at the end of $\tau=\beta/2$ evolution gives the approximation to the Gibbs state of our system Hamiltonian, $H(\vec{h})$. For more detailed calculations, please refer to Zoufal et al. [11] and McArdle et al. [33]. In this work, the technique has been adapted solely for comparative study using specific examples, with the variational quantum imaginary time evolution simulated using the built-in library available in Qiskit.
@@ B.2 Quantum Boltzmann Machine (QBM)
For a configuration |$\ket{v}$ of the visible qubits, a projective measurement on this visible state is represented as $\hat{\Pi}_{\mathbf{v}}=\ket{v}\bra{v}^{\mathbf{v}}\otimes\mathbb{I}^{\mathbf{h}}$, where $\mathbf{v}$ and $\mathbf{h}$ indicate the visible and hidden subsystems, respectively. The marginal probability of observing the visible configuration $\mathbf{v}$ is then given by,
@@
$H=-J\sigma_{1}^{x}\sigma_{2}^{x}$ $1$ $\mathbf{0.04\pm 0.01}$ $0.002\pm 0.001$
@@ Appendix C Preparing Gibbs state of a Heisenberg Hamiltonian with external field
The parameter values used for training Meta-VQT (NN Meta-VQT) are $10$ uniformly (randomly) selected values between $-2$ and $2$ for each parameter $J$ and $h$. The average trace distance between the prepared and exact Gibbs state is evaluated and compared with the VarQITE technique existing in literature as shown in Table 3.
@@ Appendix E Pseudoalgorithms
         $\tilde{\rho}\leftarrow\text{Trace out ancillas}(\text{Output state})$
         $G(\vec{h}_{i})\leftarrow\text{Evaluate Gibbs free energy}(\tilde{\rho})$
@@
[…]
@@
         Encoding neural $\text{Encoding neural network}\leftarrow(\vec{h}_{i},\vec{\phi}_{nn})$
@@
[…]
         $\tilde{\rho}\leftarrow\text{Trace out ancillas}(\text{Output state})$
         $G(\vec{h}_{i})\leftarrow\text{Evaluate Gibbs free energy}(\tilde{\rho})$
@@
[…]

### v2

@@ IV.2 Preparing Gibbs state across different regimes of Kitaev ring Hamiltonian
Here, we consider a system size of $L=3$ at temperature of $T=1/\beta=0.1$. The Meta-VQT with $6$ total qubits, $4$ encoding and $1$ processing layer is trained for the subsequent generation of quantum Gibbs state as a function of the strength of transverse field $h$. A total of 2020 $h$ values are chosen uniformly from the interval $[0.7,1.2]$ for training purposes. The reason for this particular choice of interval stems from the fact that as $h$ varies from $0.7$ to $1.2$ the model exhibits transition from semi-classical to quantum critical to semi-classical regime for a temperature value of $T=0.1$, see Fig. 10(b) and (c). For testing the Meta-VQT algorithm, $40$ uniform $h$ values are taken from the interval between $[0.7,1.2]$.
@@ IV.3 Application of NN Meta-VQT: Quantum Boltzman machine
Figure 13: NN Meta-VQT in QBM workflow : (a) and (b) shows the convergence of loss function and the trace distance between exact and prepared Gibbs state using our method of the NN Meta-VQT and existing method of the VarQITE respectively. The KL divergence in (a) converges to $0.0012$ while that in (b) converges to $0.0067$. Additionally, the prepared Gibbs state is found to be more accurate and precise for NN Meta-VQT based QBM than VarQITE based QBM. The number of layers considered for NN Meta-VQT is 44 SU2 while VarQITE requires $32$ layers of SU2 circuit to prepare Gibbs state with an average value of trace distance $0.22\pm 0.094$.
During the training phase, the trace distance between the exact and prepared Gibbs state from NN Meta-VQT and VarQITE (red triangles) is plotted in Fig. 13(a) and (b). For NN Meta-VQT based QBM, the average trace distance during the training phase is $0.071\pm 0.013$, and remains much lower than the one for VarQITE based QBM that has a value of $0.22\pm 0.094$. The circuit depth for VarQITE is $32$, which is also much larger than the depth required by NN Meta-VQT to achieve this precision. The runtime for training NN Meta-VQT is $\approx 40~\text{minutes}$ and with this trained ansatz the QBM training takes $16~\text{seconds}$. However, for VarQITE based QBM the total runtime is around $1200~\text{minutes}$ which is $30$ times more than the total runtime of NN Meta-VQT based QBM which takes $\approx 40~\text{minutes}~16~\text{seconds}$.
@@ V Conclusions
In this work, we introduced two methods of collective optimization known as Meta-VQT and NN Meta-VQT to prepare thermal states of a parametrized Hamiltonian. Unlike the traditional VQT, where the optimization is done for a single parameter configuration of the Hamiltonian, the Meta and NN Meta-VQT is trained to learn the Gibbs state of a set of parameters of the Hamiltonian. This enables efficient generalization across a distribution of Hamiltonians, reducing quantum resource overhead, and allowing rapid inference of thermal states for unseen configurations without retraining from scratch. We accomplish this by encoding the parameters of the Hamiltonian using an encoding layer, which is a SU2 hardware efficient ansatz in Meta-VQT and a neural network in NN Meta-VQT, followed by a processing layer which is HVA in Meta-VQT (Fig. 4) and hardware efficient ansatz and//or HVA in NN Meta-VQT (Fig. 5). The trainable angles in Meta-VQT is provided by the encoding and processing quantum circuits, while for NN Meta-VQT the trainable parameters are in the classical neural network. We observe that the ansatz (Meta-VQT and NN Meta-VQT) trained for a set of $10$ parameters of the Hamiltonian is able to generate Gibbs state close to the exact one corresponding to all other parameters not in the training set, but within the range of parameters chosen for training.
@@ Appendix A Variational Quantum Imaginary Time Evolution (VarQITE)
[…]
@@
[…]
@@ B.2 Quantum Boltzmann Machine (QBM)
[…]
@@
$H=-J\sigma_{1}^{x}\sigma_{2}^{x}$ 11 $\mathbf{0.04\pm 0.01}$ $0.002\pm 0.001$
@@ Appendix C Preparing Gibbs state of a Heisenberg Hamiltonian with external field
The parameter values used for training Meta-VQT (NN Meta-VQT) are $10$ uniformly (randomly) selected values between 2-2 and $2$ for each parameter $J$ and $h$. The average trace distance between the prepared and exact Gibbs state is evaluated and compared with the VarQITE technique existing in literature as shown in Table 3.
@@ Appendix E Pseudoalgorithms
   Output state $\leftarrow\text{Simulate}(\text{Ansatz})$
   $\tilde{\rho}\leftarrow\text{Trace out ancillas}(\text{Output state})$
@@
  Gradients $\leftarrow\text{Compute gradients}(\text{Global loss})$
@@
   $\text{Encoding neural network}\leftarrow(\vec{h}_{i},\vec{\phi}_{nn})$
@@
   $\text{Ansatz}\leftarrow\text{SU2}(\vec{\theta}^{(1)})+\text{HVA}(\vec{\theta}^{(2)})$
   $\text{Output state}\leftarrow\text{Simulate}(\text{Ansatz})$
   $\tilde{\rho}\leftarrow\text{Trace out ancillas}(\text{Output state})$
@@
  $\text{Global loss}\leftarrow\text{Loss}$
  $\text{Gradients}\leftarrow\text{Compute gradients}(\text{Global loss})$
@@
  $\tilde{\rho}\leftarrow$ Trace out ancilla (Output State)($\text${Output State})